# Drivetrain simulation using variational autoencoders

### v1

@@ 4.1 Data acquisition and preprossessing
Figure 1: Distribution of experimental jerk signal-data: Torque vs. electric machine speed at which the jerk signal (represented by a point) was collected. Only the stationary measurements were considered for VAE training.
@@ 4.2 Variational autoencoders as generative models for data augmentation
An empirical hyperparameter search was conducted to optimize the model architecture and the training process. Details regarding the final architecture and selected hyperparameters can be found in Table 2.
@@ 4.2.1 Unconditional VAE
Structural similarity is measured using the Structural Similarity Index Measure (SSIM), which is 0.6082, showing moderate structural alignment between the original and reconstructed spectrograms. This suggests that while the VAE captures the general structure of the data, some finer details may be lost or smoothed out. The Peak Signal-to-Noise Ratio (PSNR), at approximately 16.7421 dB, reflects the balance between noise reduction and data fidelity in the reconstructed spectrograms. Lastly, the model’s SNR output of 9.6040 dB remains close to the SNR of the original data, showing that the VAE manages to preserve the noisy characteristics of the input without introducing significant artifacts. Overall, the VAE successfully extracts core data patterns from the noisy spectrograms, providing accurate reconstructions without overly distorting the input.
Figure 4 exemplarily illustrates the performance of the unconditional VAE for a specific jerk signal. Figure 4a compares the original and generated jerk signals in the time domain, showing strong alignment, indicating good reconstruction capability. In Figure 4b, the original spectrogram is displayed, while Figure 4c shows the generated spectrogram of the VAE, which closely resembles the original. Figure 4d depicts the absolute error between the original and generated spectrograms. Local discrepancies, observed as pronounced peaks in the absolute error plot, can be attributed to noise in the original spectrograms, whereas the VAE tends to reconstruct a smoothed version. These discrepancies slightly increase the overall error, but the average error remains relatively low, as evidenced in Table 3, suggesting that the VAE has effectively learned the underlying structure of the data despite the noise, thus providing a smooth but robust reconstruction of the original jerk signal.
@@
Figure 5 displays 10,000 realizations of jerk signals generated with the trained VAE model. The original jerk signal is reconstructed from its spectrogram using an inverse STFT, where, for simplicity, the phase of the signal has been kept the same as for the original signal. The close alignment between the original jerk signal and the envelope of the generated signals demonstrates the VAE’s ability to capture the underlying pattern of the data.
Figure 5: Reparametrization trick - VAE generated jerk signals: Comparison of the original jerk signal (red) and multiple VAE-generated signals (blue shaded), showing the variability around the original signal from latent space sampling using the reparametrization trick in Equation (3.19).
Latent space interpretation and generation of new jerk signals: VAEs have an inherent limitation: they generate new outputs (spectrograms) that generally cannot be directly controlled on the basis of specific inputs (features). This drawback can be addressed by leveraging the VAE’s latent space. For instance, t-Distributed Stochastic Neighbor Embedding (t-SNE) can be used to transform the 64-dimensional latent representations into a 2D space, making it interpretable through visualizations like scatter plots [49, 50]. Then, we can generate new spectrograms $\mathbf{x}$ that reflect specific characteristics or patterns of the original data. A well-distributed latent space correlates directly with the generational capabilities of the VAE architecture. It influences the variety of features captured when sampling new data points.
@@
    First, the latent vectors are projected onto a reduced representation using the t-SNE algorithm. That is, $\mathbf{Z}=\left[\mathbf{z}_{1},...,\mathbf{z}_{n_{s}}\right]\in\mathbb{R}^{{n_{z}\times n_{s}}}$, where $n_{z}=64$ is the dimension of the latent space and $n_{s}=320$ the number of spectrograms present in the original dataset, is mapped as,
    $\mathbf{Z}_{\text{t-SNE}}=\text{t-SNE}(\mathbf{Z}),$ (4.1)
    with $\mathbf{Z}_{\text{t-SNE}}=\left[\mathbf{z}_{\text{t-SNE},1},...,\mathbf{z}_{\text{t-SNE},n_{s}}\right]\in\mathbb{R}^{{n_{\text{t-SNE}}\times n_{s}}}$. Here, $n_{\text{t-SNE}}=2$ denotes the dimension of the reduced latent space.
@@
    In the reduced space, the reduced latent vectors in $\mathbf{Z}_{\text{t-SNE}}$ are clustered and labeled according to given features of interest. Within each labeled cluster, the mean $\bm{\mu}_{\text{t-SNE}}$ and variance $\bm{{\sigma}}_{\text{t-SNE}}^{2}$ are computed. This allows us to generate new samples $\mathbf{z}_{\text{t-SNE},k}$ from a Gaussian distribution,
    ${\mathbf{z}}_{\text{t-SNE},k}\sim\mathcal{N}(\bm{{\mu}}_{\text{t-SNE}},\bm{{\sigma}}_{\text{t-SNE}}^{2}).$ (4.2)
@@
    The sampled reduced latent vectors ${\mathbf{Z}}_{\text{t-SNE}}$ are mapped back to the original latent space,
    ${\mathbf{Z}}=\text{Reconstruct}({\mathbf{Z}}_{\text{t-SNE}}).$ (4.3)
@@
    Finally, a spectrogram ${\mathbf{x}}_{k}$ is generated by passing the latent vector ${\mathbf{z}}_{k}$ through the decoder of the VAE. Specifically, the spectrograms are sampled from the likelihood $p_{\bm{\theta}}({\mathbf{x}}_{k}|\mathbf{z}_{k})$,
    ${\mathbf{x}}_{k}\sim p_{\bm{\theta}}({\mathbf{x}}_{k}|\mathbf{z}_{k}).$ (4.4)
@@
Figure 6 shows the latent space of the unconditional VAE model projected onto 2 t-SNE components, where the data points are categorized by a vehicle type and b predefined torque ranges. The data is labeled according to the two vehicle types (Figure 6a) and seven torque ranges (Figure 6b), which are subsets of the interval $\left[-300,1000\right]$ Nm. A different marker shape represents each torque range, and ellipses are drawn around clusters based on their covariance to highlight the underlying distribution. It is observed that vehicle type better categorizes the data than torque range. However, lower (0 - 50 Nm) and higher (600 - 1000 Nm) torque categories show tighter clustering compared to other ranges, as seen by their compact ellipses. The other torque categories tend to overlap more, suggesting that the latent representations for these torque ranges may not be as easily distinguished from each other. This overlap is common when the VAE finds similarities in the latent features of different categories. Overall, the plot highlights the distribution and spread of the data in the latent space, and the presence of distinct clusters suggests that the VAE has learned meaningful latent features, particularly with respect to vehicle types.
@@
The spectrograms in Figure 7a display frequency components over time for the generated jerk signals, demonstrating how the VAE can reconstruct plausible frequency representations. Figure 7b shows the jerk signals generated in the time domain, where variations in the torque range lead to different signal shapes. Figure 7c shows the frequency spectrum of each jerk signal. The jerk signals in the training and testing datasets used for VAE training consistently exhibit their highest amplitudes around 1 Hz and 10 Hz. Notice that these frequencies align with vibrations in the drivetrain’s natural frequency range caused by abrupt acceleration changes. Therefore, we shaded the frequency bands (0–2 Hz and 8–12 Hz) to emphasize that the generated signals are physically plausible. Griffin-Lim algorithm [40] is applied here to estimate the phase (see Section 3.2) by starting with a random initialization and iteratively refining the phase estimate. An accurate estimation of the phase was achieved after $1,000$ iterations, which took less than 1 second to complete.
@@ 4.2.2 Conditional VAEs
Performance evaluation: Table 3 compares the unconditional VAE, CVAE, and GMM-CVAE performance. The table reveals that the unconditional VAE consistently performs slightly better on most metrics. It achieves the lowest reconstruction errors (MSE = 0.3654, MAE = 0.44291), suggesting that it is more effective in accurately reconstructing noisy spectrograms. The unconditional VAE also shows a higher structural similarity (SSIM = 0.6082) compared to CVAE (SSIM = 0.6051) and GMM-CVAE (SSIM = 0.5759), indicating that it preserves more structural information in the data. Moreover, it outperforms the other models in PSNR (16.7421 dB) and SNR (9.6040 dB), suggesting that it reduces noise more effectively while maintaining signal fidelity. Although CVAE shows comparable performance, it does not provide significant improvements in terms of accuracy. The GMM-CVAE shows higher reconstruction errors and slightly lower SSIM, PSNR, and SNR, indicating that the additional complexity does not result in better performance in this case. Overall, the unconditional VAE proves to be the most effective model on the testing set, capturing the core features of the data while handling noise more effectively than the other models. However, the CVAEs are better suited for the generation of new jerk signals given a desired feature (i.e., condition), such as a torque demand.
@@
As in the unconditional case, Figures 8 and A3 show that the conditional VAEs have effectively learned the underlying distribution of the jerk signals. Unlike the unconditional VAE, which could only generate jerk signals from labeled torque ranges, the conditional VAE allows for generating jerk signals by specifying a precise torque value as a condition. This provides a significant advantage when using the VAE as a replacement in drivetrain simulations, where control over specific input parameters is a desired feature. The generated signals are also physically plausible in the CVAE and GMM-CVAE cases, as the jerk signals in the dataset used for training correspond to frequency regions with the highest amplitudes around 1 and 10 Hz.
@@ 4.3 Comparison between conditional generative-, physics-, and hybrid-based models
In this section, we compare the predictive performance of a physics-based drivetrain (hardware) model, a hybrid model, and the CVAE generative model for the dataset presented in Section 4.1. The CVAE was chosen here because it represents the best trade-off between accuracy and versatility among the explored VAE architectures; see Table 3. In particular, the unconditional VAE has the limitation that the generated jerk signal cannot be conditioned to a specific torque value, while the GMM-CVAE does not provide a significant improvement in performance over the CVAE and introduces additional complexity in sampling from the approximated likelihood. We refer the reader to Appendix C for details of the setup of physics-based and hybrid models.
Table 4 summarizes the MSE values for the jerk signals generated by each model. As the hardware model is a simple spring-mass oscillator system, it results in a much lower accuracy than other models. This could be improved by using a more sophisticated hardware model, which was, however, out of scope for this study. Also note that our goal is to increase the efficiency of the drivetrain development process, and the development of such hardware models is generally time-consuming.
The hybrid model, which utilizes a 7-million-parameter U-Net convolutional neural network architecture (detailed in Appendix C), demonstrates slightly better accuracy than the CVAE. However, its substantially higher number of parameters makes it significantly more data intensive and challenging to train compared to the VAEs tested in this study. This complexity underscores the trade-off between accuracy and model efficiency, with VAEs offering a more practical alternative for scenarios with limited data availability.
Metric Hardware Model Hybrid Model CVAE
@@
Table 4: Comparison physics-based, hybrid, and CVAE as a generative model. MSE values for hardware simulator model, hybrid model, and CVAE. The MSE was computed as the mean squared error between the actual and generated signals (not spectrograms) over the test set. Results were averaged across the test set size to ensure comparability between models. For instance, for the hybrid model, the MSE was calculated over 238 jerk samples.

### v2

@@ 4.1 Data acquisition and preprossessing
Figure 1: Distribution of the available experimental jerk signal-data: Torque vs. electric machine speed at which the jerk signal (represented by a point) was collected. Only the stationary measurements derived from these signals were considered for VAE training. Notice that the observed non-uniformity in the sampling reflects the inherent distribution within the collected measurement campaign, not a specific sampling choice for this visualization.
@@
The reader should notice that the time-domain signals and their corresponding spectrograms may exhibit a discrepancy in their time axes. This is due to the nature of the STFT. The spectrogram represents overlapping windowed segments of the signal, where the time resolution depends on the hop size and window function specified in Table 1. A small hop size, in this case $2$, results in a higher number of time frames, stretching the spectrogram along the time axis. This discrepancy is a consequence of time-frequency analysis and does not indicate a misalignment in data representation.
@@ 4.2 Variational autoencoders as generative models for data augmentation
An empirical hyperparameter search was conducted to optimize the model architecture (including aspects like the number of layers, filter sizes, and latent dimension) and the training process (such as learning rate and choice of optimizer). This involved iterative experimentation and evaluation on a validation set, guided by common practices in VAE applications. Details regarding the final architecture and selected hyperparameters can be found in Table 2.
@@ 4.2.1 Unconditional VAE
Structural similarity is measured using the Structural Similarity Index Measure (SSIM), which is 0.6082, showing moderate structural alignment between the original and reconstructed spectrograms. This suggests that while the VAE captures the general structure of the data, some finer details may be lost or smoothed out. The Peak Signal-to-Noise Ratio (PSNR), at approximately 16.7421 dB for the unconditional VAE, reflects the balance between noise reduction and data fidelity in the reconstructed spectrograms. While this PSNR value might be considered moderate in some image reconstruction tasks, it is important to interpret it in the context of VAEs and the nature of our data. VAEs are designed to learn a compressed and regularized latent representation, which often results in a smoothing or denoising effect on the reconstructions, particularly for inherently noisy input data (our original spectrograms have an SNR of about 5.5 dB). This smoothing can result in the loss of some fine-grained details or noise present in the original, thereby reducing the PSNR. This is a consequence of the model prioritizing a generalizable latent space and capturing the underlying data structure over achieving perfect pixel-wise reconstruction. Lastly, the model’s SNR output of 9.6040 dB indicates an improvement over the original data’s SNR, suggesting the VAE does achieve some denoising while preserving essential signal characteristics. Overall, the VAE successfully extracts core data patterns from the noisy spectrograms, providing accurate reconstructions without overly distorting the input.
Figure 4 exemplarily illustrates the performance of the unconditional VAE for a specific jerk signal. Figure 4a compares the original and generated jerk signals in the time domain, showing strong alignment, indicating good reconstruction capability. In Figure 4b, the original spectrogram is displayed, while Figure 4c shows the generated spectrogram of the VAE, which closely resembles the original. Figure 4d depicts the absolute error between the original and generated spectrograms. Local discrepancies, observed as pronounced peaks in the absolute error plot, can be attributed to noise in the original spectrograms. As characteristic of VAEs, particularly due to the regularization imposed by the KL divergence term and the bottleneck structure, the model tends to reconstruct a smoothed version of the input, effectively filtering some of the high-frequency noise. This behavior is evident in Figure 4c appearing less noisy than Figure 4b. These discrepancies and the loss of some noise components contribute to the reconstruction error but also highlight the VAE’s ability to capture the more salient underlying features. The average error remains relatively low (Table 3), suggesting that the VAE has effectively learned this underlying structure, providing a smooth yet robust reconstruction of the original jerk signal.
@@
Figure 5 displays 10,000 realizations of jerk signals. These signals are generated by first encoding the original jerk signal (red) to obtain its latent distribution $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}_{\text{original}})$, then repeatedly sampling $\mathbf{z}$ from this distribution using the reparameterization trick (Equation 3.19), and finally decoding these latent samples. The phase for all signals is estimated using the Griffin-Lim algorithm. The close alignment between the original jerk signal and the envelope of these generated signals demonstrates the VAE’s ability to learn a consistent encoding and to represent local variability around a given input.
Figure 5: Reparametrization trick - VAE generated jerk signal variability: Comparison of the original jerk signal (red) with the mean (e.g., solid blue line) and standard deviation envelope (e.g., shaded blue region) of 10,000 jerk signals generated by sampling from the latent posterior $q_{\bm{\phi}}(\mathbf{z}|\mathbf{x}_{\text{original}})$ using the reparameterization trick (Equation 3.19) and then decoding. The decoder output variance $\lambda$ is fixed (see Section 3.2), so the displayed variability originates from the latent space sampling.
Latent space interpretation and generation of new jerk signals: VAEs have an inherent limitation: they generate new outputs (spectrograms) that generally cannot be directly controlled on the basis of specific inputs (features). This drawback can be addressed by leveraging the VAE’s latent space. For instance, t-Distributed Stochastic Neighbor Embedding (t-SNE) can be used to transform the 64-dimensional latent representations into a 2D space, making it interpretable through visualizations like scatter plots [49, 50]. Then, we can generate new spectrograms $\mathbf{x}$ that reflect specific characteristics or patterns of the original data. A well-structured (or "well-distributed") latent space correlates directly with the generative capabilities of the VAE architecture. Such a latent space ideally exhibits desirable properties like continuity (whereby small changes in the latent vector lead to correspondingly small changes in the generated output), smoothness, and a meaningful organization where different data variations are encoded in a somewhat disentangled or structured manner. These characteristics are crucial for enabling robust interpolation between data points and controlled generation of novel samples. It influences the variety of features captured when sampling new data points and the overall quality and coherence of the generated data.
@@
    The set of original latent vectors $\mathbf{Z}_{train}=\{\mathbf{z}^{(i)}\}_{i=1}^{N_{train}}\subset\mathbb{R}^{n_{z}}$ obtained by encoding the training dataset spectrograms is projected into a 2D space using the t-SNE algorithm:
    $\mathbf{Z}_{\text{t-SNE},train}=\text{t-SNE}(\mathbf{Z}_{train}),$ (4.1)
    where $\mathbf{Z}_{\text{t-SNE},train}=\{\mathbf{z}^{(i)}_{\text{t-SNE}}\}_{i=1}^{N_{train}}\subset\mathbb{R}^{2}$. This 2D embedding is used for visualization.
@@
    For a chosen pre-defined category of interest (e.g., a specific torque range), we identify the subset of training data points belonging to this category. Let this subset of 2D t-SNE embeddings be $\mathbf{Z}^{\prime}_{\text{t-SNE}}\subset\mathbf{Z}_{\text{t-SNE},train}$. We then compute the empirical mean $\bm{\mu}^{\prime}_{\text{t-SNE}}$ and covariance $\bm{\Sigma}^{\prime}_{\text{t-SNE}}$ of the points in $\mathbf{Z}^{\prime}_{\text{t-SNE}}$. New points in the 2D visualization space, denoted $\mathbf{z}_{\text{t-SNE},new}$, are sampled from a Gaussian distribution parameterized by these statistics:
    $\mathbf{z}_{\text{t-SNE},new}\sim\mathcal{N}(\bm{\mu}^{\prime}_{\text{t-SNE}},\bm{\Sigma}^{\prime}_{\text{t-SNE}}).$ (4.2)
@@
    To obtain a corresponding high-dimensional latent vector $\mathbf{z}_{new}\in\mathbb{R}^{n_{z}}$ for each $\mathbf{z}_{\text{t-SNE},new}$, we employ the k-Nearest Neighbors (k-NN) algorithm available in the Python library scikit-learn [51, 52]. Specifically, for a given $\mathbf{z}_{\text{t-SNE},new}$, we identify the $k$ training data points in $\mathbf{Z}_{\text{t-SNE},train}$ that are closest (in Euclidean distance) to $\mathbf{z}_{\text{t-SNE},new}$. Let their corresponding original high-dimensional latent vectors be $\{\mathbf{z}^{*(1)},...,\mathbf{z}^{*(k)}\}\subset\mathbf{Z}_{train}$. The new high-dimensional latent vector $\mathbf{z}_{new}$ is then computed as a weighted average of these $k$ neighbors:
    $\mathbf{z}_{new}=\sum_{j=1}^{k}w_{j}\mathbf{z}^{*(j)},\quad\text{where }w_{j}=\frac{\,\text{exp}{\left(-d(\mathbf{z}_{\text{t-SNE},new},\mathbf{z}^{(j)}_{\text{t-SNE}})/\tau\right)}}{\sum_{l=1}^{k}\,\text{exp}{\left(-d(\mathbf{z}_{\text{t-SNE},new},\mathbf{z}^{(l)}_{\text{t-SNE}})/\tau\right)}},$ (4.3)
    $d(\cdot,\cdot)$ is the Euclidean distance, and $\tau$ is a parameter controlling the smoothness (larger $\tau$) or sharpness (smaller $\tau$) of the weighting distribution applied to the k-NN. In our implementation, we used $k=5$ and $\tau=0.1$. This step aims to find a plausible point in the original latent space that is consistent with the location sampled in the 2D visualization.
@@
    Finally, a new spectrogram $\mathbf{x}_{new}$ is generated by passing the derived high-dimensional latent vector $\mathbf{z}_{new}$ through the VAE’s decoder:
    $\mathbf{x}_{new}\sim p_{\bm{\theta}}(\mathbf{x}\mid\mathbf{z}_{new}).$ (4.4)
This procedure leverages the t-SNE visualization to identify regions associated with known physical characteristics (e.g., specific torque ranges). By sampling from these regions in the 2D t-SNE space and then mapping back to the original latent space, the aim is to generate new data that shares these desired physical attributes. However, it is important to acknowledge that t-SNE is a non-linear dimensionality reduction technique primarily designed for visualization, and it may not preserve the global structure of the data perfectly. Therefore, the process of sampling in the t-SNE space and then reconstructing to the original latent space is an exploratory method with inherent limitations regarding the precision of the inverse mapping. The resulting samples’ appropriateness for the target physical category depends on how well the VAE’s latent space is organized.
@@
Figure 6 shows the latent space of the unconditional VAE model, obtained by mapping the 64-dimensional latent vectors $\mathbf{Z}$ to a 2D representation $\mathbf{Z}_{\text{t-SNE}}$ using t-SNE for visualization purposes. The data points in this 2D projection are colored and shaped (i.e., labeled) by the two vehicle types (Figure 6a) and seven torque ranges (Figure 6b), which are subsets of the interval $\left[-300,1000\right]$ Nm. A different marker shape represents each torque range, and ellipses are drawn around these predefined groups based on their covariance in the 2D t-SNE space to highlight their approximate distribution and separation. It is observed that vehicle type better categorizes the data than torque range. However, lower (0 - 50 Nm) and higher (600 - 1000 Nm) torque categories show tighter clustering compared to other ranges, as seen by their compact ellipses. The other torque categories tend to overlap more, suggesting that the latent representations for these torque ranges may not be as easily distinguished from each other. This overlap is common when the VAE finds similarities in the latent features of different categories. Overall, the plot highlights the distribution and spread of the data in the latent space, and the presence of distinct clusters suggests that the VAE has learned meaningful latent features, particularly with respect to vehicle types.
@@
The spectrograms in Figure 7a display frequency components over time for the generated jerk signals, demonstrating how the VAE can reconstruct plausible frequency representations. Figure 7b shows the jerk signals generated in the time domain, where variations in the torque range lead to different signal shapes. Figure 7c shows the frequency spectrum of each jerk signal. The jerk signals in the training and testing datasets used for VAE training consistently exhibit their highest amplitudes around 0-2 Hz and 8-12 Hz. Notice that these frequencies align with vibrations in the drivetrain’s natural frequency range caused by abrupt acceleration changes. Therefore, we shaded the frequency bands (0–2 Hz and 8–12 Hz) to emphasize that the generated signals are physically plausible. Griffin-Lim algorithm [40] is applied here to estimate the phase (see Section 3.2) by starting with a random initialization and iteratively refining the phase estimate. An accurate estimation of the phase was achieved after $1,000$ iterations, which took less than 1 second to complete.
@@ 4.2.2 Conditional VAEs
Performance evaluation: Table 3 compares the unconditional VAE, CVAE, and GMM-CVAE performance. The table reveals that the unconditional VAE consistently performs slightly better on most metrics when evaluated on spectrogram reconstruction. It achieves the lowest reconstruction errors (MSE = 0.3551, MAE = 0.4203), suggesting that it is more effective in accurately reconstructing noisy spectrograms. The unconditional VAE also shows a higher structural similarity (SSIM = 0.6082) compared to CVAE (SSIM = 0.6051) and GMM-CVAE (SSIM = 0.5759), indicating that it preserves more structural information in the data. Moreover, it outperforms the other models in PSNR (16.7421 dB) and SNR (9.6040 dB), suggesting that it reduces noise more effectively while maintaining signal fidelity. Although CVAE shows comparable performance, it does not provide significant improvements in terms of accuracy. The GMM-CVAE shows higher reconstruction errors and slightly lower SSIM, PSNR, and SNR, indicating that the additional complexity does not result in better performance in this case. Overall, the unconditional VAE proves to be the most effective model on the testing set for spectrogram reconstruction, capturing the core features of the data while handling noise more effectively than the other VAE variants. While the differences in these spectrogram-domain metrics (like PSNR) between the VAE variants are relatively small, these metrics reflect the direct output of the models. A comparison of generated time-domain jerk signals, also against other modeling approaches, is provided in Section 4.3. However, the CVAEs offer the advantage of conditional generation for specific features like torque demand, which is a key capability for targeted simulations.
The performance metrics for the VAE variants reported in Table 3 are based on a single, fixed random split of the 320 stationary spectrograms into training, validation, and a test set of 32 samples. While this provides a consistent basis for comparison among the VAE models, it is acknowledged that with a test set of this size, the precise metric values might exhibit some sensitivity to different data splits; an exhaustive analysis of this sensitivity (e.g., via cross-validation) was not performed in the current study.
@@
As in the unconditional case, Figures 8 and A3 show that the conditional VAEs have effectively learned the underlying distribution of the jerk signals. Unlike the unconditional VAE, which could only generate jerk signals from labeled torque ranges, the conditional VAE allows for generating jerk signals by specifying a precise torque value as a condition. This provides a significant advantage when using the VAE as a replacement in drivetrain simulations, where control over specific input parameters is a desired feature. The generated signals are also physically plausible in the CVAE and GMM-CVAE cases, as the jerk signals in the dataset used for training correspond to frequency regions with the highest amplitudes around 0-2 Hz and 8-12 Hz.
When comparing the jerk signals generated by the CVAE (e.g., Figure 8) with those from the unconditional VAE for similar torque ranges (e.g., Figure 7), some differences in character may be observed. This can be attributed to the distinct ways the latent spaces are structured and utilized. The unconditional VAE learns a general latent representation, and sampling from visualized torque-range regions taps into this shared space. In contrast, the CVAE’s latent space is explicitly conditioned on the torque demand. This conditioning tends to create a more structured latent space where specific torque inputs map to more defined regions, leading to generated signals that are highly consistent with the given condition, potentially with different diversity characteristics compared to samples from the unconditioned VAE for a broader torque category.
@@ 4.3 Comparison between conditional generative-, physics-, and hybrid-based models
In this section, we compare the predictive performance of a physics-based drivetrain model (referred to as physics-based model), a hybrid model, and the CVAE generative model for the dataset presented in Section 4.1. The CVAE was chosen here because it represents the best trade-off between accuracy and versatility among the explored VAE architectures; see Table 3. In particular, the unconditional VAE has the limitation that the generated jerk signal cannot be conditioned to a specific torque value, while the GMM-CVAE does not provide a significant improvement in performance over the CVAE and introduces additional complexity in sampling from the approximated likelihood. We refer the reader to Appendix C for details of the setup of physics-based and hybrid models.
Table 4 summarizes the MSE values for the jerk signals generated by each model. As the physics-based model is a simplified representation (specifically, the two-mass spring-mass oscillator system mentioned in Section 2), it results in a much lower accuracy compared to the data-driven and hybrid approaches. This simplified model serves as a basic benchmark and is not intended to be representative of the sophisticated, highly calibrated physics-based simulation tools used in detailed industrial applications. Its performance underscores the difficulty in achieving high fidelity with basic physical formulations alone and motivates the use of data-driven techniques like VAEs or hybrid models to capture more complex system behaviors. This could be improved by using a more sophisticated physics-based model, which was, however, out of scope for this study. Also note that our goal is to increase the efficiency of the drivetrain development process, and the development of such physics-based models is generally time-consuming.
The hybrid model, which utilizes a 7-million-parameter U-Net convolutional neural network architecture as a data-driven correction stage on top of a basic drivetrain physics simulator (see Appendix C for details), demonstrates slightly better accuracy in predicting time-domain jerk signals than the CVAE when the CVAE is used generatively for a given condition (Table 4). We acknowledge that the U-Net in the hybrid model has significantly more parameters (7 million) compared to our VAE architectures. This notable difference in model scale (roughly a factor of 15-20) means the comparison is not strictly like-for-like in terms of model complexity or data requirements, although the discrepancy is not as large as three orders of magnitude. The primary focus of the VAE development was on generative capabilities and interoperability with more compact models. The U-Net hybrid model is presented here as a benchmark, representing a high-performance approach for signal prediction in this domain, which helps contextualize the CVAE’s generative performance. Its substantially higher number of parameters makes it significantly more data-intensive and challenging to train compared to the VAEs tested in this study. This complexity underscores the trade-off between accuracy and model efficiency (referring to aspects like lower data requirements for training, fewer model parameters, and potentially faster inference or training times), with VAEs offering a more practical alternative for scenarios with limited data availability or computational resources. A detailed comparison with hybrid models having a parameter count similar to the VAEs was not performed in this study but could be a direction for future work to assess parameter efficiency more directly for the predictive task.
Metric Physics-based Model Hybrid Model CVAE
@@
Table 4: Comparison of physics-based model, hybrid, and CVAE as a generative model. The MSE was computed as the mean squared error between the actual and generated signals (not spectrograms) over the common test set consisting of 32 jerk signals (derived from the stationary signals described in Section 4.1). All models in this table were evaluated on these same 32 test signals to ensure comparability.